Hercules Films is deciding on the price of the video release of its film Son of Frankenstein. Its marketing people estimate that at a price of p dollars, it can sell a total of q = 200,000-20,000p copies. What price will bring in the greatest revenue?
Answer by ikleyn(52778)

.
The revenue = the product of the price by the quantity sold:  R = p*q,   or


R = p*(200000-20000*p) = 20000*p*(10-p).


The roots of this quadratic function are p= 0  and  p= 10;  hence, the maximum of the quadratic function is located 
exactly half way between the roots, i.e. at p= 5.


Answer.  The price p= 5 dollars per video provides the maximum revenue.